# Experimentally Testing AI-Powered Content Transformations on Student Learning

### v1

@@
Abstract.
Generative AI, which is capable of transforming static content into dynamic learning experiences, holds the potential to revolutionize student engagement in educational contexts. However, questions still remain around whether or not these tools are effective at facilitating student learning. In this research, we test the effectiveness of content transformations through Learn Your Way, an experimental research platform that transforms textbook chapters into dynamic visual and audio representations. Through a between-subjects, mixed methods experiment with 60 US-based students, we demonstrate that students who used Learn Your Way had a more positive learning experience and had better learning outcomes compared to students learning the same content via a digital reader. These findings indicate that AI-driven tools, capable of providing multiple, interactive representations of content, constitute an effective and promising method for enhancing student learning.
@@ Abstract.
copyright: noneccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing User studiesccs: Human-centered computing Laboratory experiments
1. Introduction
@@
Around the world, AI is being leveraged to transform lives and augment human capabilities in multiple industries, and education is no exception. Teachers, students, and other stakeholders in the industry have been known to leverage AI to support lesson planning and delivery, learning, administrative tasks, etc. (Łuczak et al., 2024).
@@
With the continuous push towards digital equity and literacy over the years, there has been an increased interest in education technologies (particularly chat-based tools like ChatGPT [^1] https://chatgpt.com/ and Gemini [^2] https://gemini.google.com/) (Zhai et al., 2021; Zhang and Aslan, 2021; Adeshola and Adepoju, 2024; Memarian and Doleck, 2023; Li et al., 2025). Amongst other benefits, generative AI is uniquely poised to transform learning materials (including traditional textbooks) into more dynamic, engaging, and personalized content and to provide immediate feedback to improve understanding of learning materials (Łuczak et al., 2024).
@@
As improvements are made in natural language processing and understanding, large language models (LLMs) show promise for generating effective learning resources at scale for diverse groups of learners. This goes beyond performance outcomes to other intrinsic factors such as interest and motivation in learning. For example, in a between-subjects study comparing pre-existing text to AI generated output, Leong et al. show that while there was no difference in learning performance, AI generated output positively affected learning motivation thereby facilitating positive downstream effects (including appreciating the relevance and utility of learning the material) (Leong et al., 2024).
@@
Research in education also shows that learners benefit from having information presented in multiple formats (Mayer, 2002). By presenting information in multiple modalities and pairing these transformations with guided activities, reflection, and feedback, learners benefit from numerous pedagogical advantages, including the promotion of independent learning and practice that is tailored to the needs of individual students (Moreno and Mayer, 2007).
@@
The process of transforming traditional textbooks and learning materials into these more dynamic, engaging, and personalized sources can, however, be time consuming and labor intensive for teachers, thereby limiting its broader use (Dennison et al., 2025). When considering direct and independent use by students for more personalized experiences at scale, several areas of concern emerge that demand careful consideration. These include the critical issues of accuracy, bias, academic integrity, factuality and reliability of model outputs, the potential for over-reliance on AI tools and its long-term impact on cognitive abilities and critical thinking skills, and students’ inherent capacity to formulate effective prompts for facilitating their learning processes (Han et al., 2024; Park and Ahn, 2024; Tankelevitch et al., 2025; Chan and Lee, 2023; Denny et al., 2023). Furthermore, research has been mixed when it comes to demonstrating the efficacy of AI tools on performance outcomes (Do et al., 2025; Leong et al., 2024; Pallant et al., 2025).
@@
In this paper, we build upon pedagogical best practices to experimentally test the impact of AI-generated content transformations and feedback on student understanding and recall of information. We investigate the following research questions:
  • RQ1:
@@
    How does Learn Your Way, an experimental research platform that leverages generative-AI to transform content (Wang et al., 2025), impact student comprehension and immediate recall of information?
  • RQ2:
@@
    How does Learn Your Way impact retention and recall of information over time?
  • RQ3:
@@
    Can Learn Your Way facilitate more engaging learning experiences for students?
@@
The results of the experiment suggest that AI-powered content transformations can facilitate positive learning outcomes and provide students with more engaging learning experiences compared to learning via traditional textbook chapters. We also demonstrate the role of timely feedback (as opposed to solutions) in equipping learners with knowledge and skills for self efficacy and academic independence.
2. Background and Related Work
2.1. HCI and Education
@@
The Human-Computer Interaction (HCI) community has a rich history of research exploring the design, development, and evaluation of educational technology (Stasko et al., 1993; Russell et al., 1993; Abowd et al., 1998; Scaife et al., 1997; Zuckerman et al., 2005; Saerbeck et al., 2010; Winkler et al., 2020). The discourse about generative AI in education has shifted from an initial period of denial and uncertainty to partial acceptance and explorations, ultimately leading to calls for further research (Pallant et al., 2025) and governance (UNICEF Innocenti, 2019). Recent HCI research has explored a wide range of important topics in this domain, including privacy and security considerations for technology use in schools (Kumar et al., 2019; Chanenson et al., 2023), extending participatory and co-design methods to the development of classroom-based technologies (Chang et al., 2024; Nicholson et al., 2022), and designing tactile-based learning experiences to foster accessibility in education (Melfi et al., 2020).
@@
Additionally, research in HCI is beginning to explore how the increase in AI-based capabilities can be incorporated into a variety of interactive experiences in order to positively impact learning. For example, Liu et al. created and evaluated an AI agent designed to interact with students in VR-based learning environments and promote classroom participation and more positive learning experiences, finding that these agents can provide pedagogical value through increased levels of engagement and the potential for learning gains (Liu et al., 2024). Chen et al. found that augmenting block-based programing software (Scratch [^3] https://scratch.mit.edu) with an LLM-based chat experience allowed students to learn more independently, and create higher quality projects (Chen et al., 2024). This growing body of research suggests a strong desire from the HCI community to better understand how AI-based tools can help improve student outcomes.
2.2. AI Assessment and Feedback
@@
One of the ways AI can help bring value to students is through AI-driven assessment and feedback. Pedagogical research shows that effective feedback is timely, actionable, grounded in standards or learning goals, and cites specific observations to facilitate learner understanding and support action (Wiggins, 2012). Providing feedback is an inherent part of a teacher’s role, but as students partake in independent practice outside of classroom hours, feedback is not always readily available or accessible. However, the ability to access feedback is necessary to help students understand their performance and progress in relation to required standards and learning objectives (Hattie and Timperley, 2007).
@@
From the early days of intelligent tutoring systems, AI-driven knowledge assessments – including personalized, adaptive feedback – has been a core area of research for education-focused AI (Anderson et al., 1985; Corbett et al., 1997; Graesser et al., 2004). With generative AI, students have the potential to access ongoing, targeted, timely, and individualized assessments and feedback at scale for self-directed learning and self-efficacy. Recent technological advances show improvements with the capabilities surrounding assessment and feedback (Choi and McClenen, 2020; Yang et al., 2022), but more improvements are needed to truly reach this vision (Tu et al., 2025).
@@
In HCI, several qualitative studies with parents, teachers, and students have led to design recommendations that emphasize the value, importance, and opportunity for AI to provide students with high-quality assessments and feedback (Han et al., 2024; Prasad et al., 2025). A recent meta-analysis on emerging technologies in education also emphasizes the importance of HCI research that is grounded in pedagogical strategies and focuses on assessment and feedback (Van Mechelen et al., 2023). However, there is limited quantitative data and experimental studies evaluating how the design of AI-based experiences can facilitate effective assessments and feedback. In this paper, we build upon prior work, by incorporating design recommendations from past research into Learn Your Way, and by assessing the impact of the system on learning performance and experience.
2.3. AI-Generated Content Transformations
@@
By definition, content transformations transform static content into multiple representations to help facilitate learning, such as transforming text-based content into something that is audio- and/or visual-based (Ainsworth, 1999). In contemporary educational contexts, there is a notable rise in the adoption of content transformation capabilities, which are emerging as a key complement to more traditional pedagogical tools like assessment and feedback. This trend reflects a broader shift towards dynamic, personalized learning resources that can be adapted to various student needs and learning preferences. The benefits of content transformations are guided by principles from Mayer’s cognitive theory of multimedia learning, which states that people learn more deeply when information is received in multiple modalities as opposed to reading or hearing words alone (Mayer, 2002; Mayer and Moreno, 1998). For example, a study showed that children who listened to stories while also viewing relevant pictures recalled more information than those who only heard stories (Levin and Berry, 1980). This study (and others, e.g., (Villegas-Ch et al., 2024)) suggests that integrating materials across different modalities, such as audio paired with images or videos, can be particularly beneficial for student engagement and learning.
@@
Generative AI capabilities, by their nature, create content transformations (text-to-text, text-to-audio, text-to-image, and text-to-video). By providing multiple modes of interaction and engagement with content, generative AI tools are well positioned to support self-directed learning, including theories that advocate for learners to play a central role in their own learning journeys (Deci and Ryan, 2000; Zimmerman and Moylan, 2009; Zimmerman, 2000). This could include metacognitive strategies that allow for setting individual learning goals, choosing appropriate strategies to achieve these goals, and monitoring one’s individual learning process and outcomes for adjustment and improvements (Panadero, 2017).
@@
In HCI, the research community is beginning to explore the effects of AI-based content transformations on learning and education. In an experiment, Leong et al. found that a generative AI-based vocabulary app that transformed standard vocabulary sentences into either personalized, generated sentences or personalized, generated text-based stories did not increase learning performance but did positively impact learning motivation (Leong et al., 2024). Similarly, in an experiment moving beyond text-to-text transformations, Do et al. found that transforming educational materials from text to podcast form led to a more enjoyable learning experience and also, significantly improved learning outcomes for certain subjects (Do et al., 2025).
@@
3. Learn Your Way
Learn Your Way is an experimental platform that leverages generative AI to transform textbook PDFs into a variety of formats, or multiple representations, previously described at (Wang et al., 2025). We used Learn Your Way to test our research hypotheses. These content transformations were grounded in pedagogical theories that demonstrate that multiple representations of content can be a powerful way to learn (Mayer and Gallini, 1990; Mayer and Moreno, 1998). The student is able to choose the format they want to engage with based on their learning preferences.
Learn Your Way presents five learning modalities: Immersive Text, Slides, Video, Audio Lesson, and Mindmap (illustrated in Figure 1). See Appendix A for a visual example of each modality.
  • Immersive Text: Transforms the original content into smaller sections releveled to a selected grade, and provides opportunities for interactive learning and practice (e.g. memory aids, quizzes, embedded questions, and timelines). Within Immersive Text, students had access to the following features:
@@ 3. Learn Your Way
Figure 1. Diagram of the Learn Your Way user interface for Immersive Text. Other learning modalities are featured at the top including Slides, Video, Audio Lesson, and Mindmap, as well as the original PDF.:
When students open Learn Your Way, they start by selecting their grade level and interests (see Appendix B). Learn Your Way then personalizes and adapts the content being generated by adjusting for grade level and incorporating the student’s selected interests. As students use Learn Your Way, they engage with interactive quizzes designed to promote active learning and assess real-time understanding of the source material, enabling further elements of personalization. Students receive dynamic feedback on their answers, can track their overall progress, and are guided to review specific content areas where they demonstrated gaps in understanding. Students are also able to view the original content, in PDF form, at any time.
3.1. Learn Your Way Design and Evaluation
Throughout the design of Learn Your Way, each modality was evaluated by pedagogical experts to ensure the fidelity and quality of each content transformation while preserving the integrity of the source content. A unified rubric was developed to assess key pedagogical criteria, including accuracy (the extent to which the AI faithfully represented the source material), coverage (the thoroughness with which all key concepts were addressed), and learning science principles. The results indicated a high level of quality, with all modalities and features achieving an average expert rating of 0.9 or higher on a scale of 0 to 1 across all pedagogical criteria. Detailed information about the rubric and results of the evaluation can be found in Appendix C.
4. Experimental Study
We conducted an experimental study to compare performance outcomes and learning experiences between the AI-powered content transformation tool (Learn Your Way) and a traditional digital reading tool (Digital Reader). The digital reader used was Adobe Acrobat Reader v25.001.20531, which is designed for viewing, printing, and signing PDF documents. It also has built-in annotation features such as highlighting and commenting. To ensure the digital reader operated as a true control condition, we turned off the AI Assistant feature (Adobe, 2024) to simulate the experience of reading a traditional textbook chapter.
The source content used in the experiment was a neuroscience textbook chapter titled, “Brain Development During Adolescence” (Learning, 2019). The source material was chosen by a panel of pedagogy experts using a set of criteria based on education psychology (see Supplementary Materials).
4.1. Research Hypotheses

### v2

@@
Abstract
Generative AI, which is capable of transforming static content into dynamic learning experiences, holds the potential to revolutionize student engagement in educational contexts. However, questions still remain around whether or not these tools are effective at facilitating student learning. In this research, we test the effectiveness of an AI-powered platform incorporating multiple representations and assessment through Learn Your Way, an experimental research platform that transforms textbook chapters into dynamic visual and audio representations. Through a between-subjects, mixed methods experiment with 60 US-based students, we demonstrate that students who used Learn Your Way had a more positive learning experience and had better learning outcomes compared to students learning the same content through a digital textbook. These findings indicate that AI-driven tools, capable of providing choice among interactive representations of content, constitute an effective and promising method for enhancing student learning.
[…]
1 Keywords:
@@
2 Introduction
@@
Around the world, AI is being leveraged to transform lives and augment human capabilities in multiple industries, and education is no exception. Teachers, students, and other stakeholders in the industry have been known to leverage AI to support lesson planning and delivery, learning, administrative tasks, etc. (LUCZAK2024).
@@
With the continuous push towards digital equity and literacy over the years, there has been an increased interest in education technologies (particularly chat-based tools like ChatGPT [^1] https://chatgpt.com/ and Gemini [^2] https://gemini.google.com/) (Zhai2021; ZHANG2021; Adeshola2024; MEMARIAN2023; Li2025). Amongst other benefits, generative AI is uniquely poised to transform learning materials (including traditional textbooks) into more dynamic, engaging, and personalized content and to provide immediate feedback to improve understanding of learning materials (LUCZAK2024).
@@
As improvements are made in natural language processing and understanding, large language models (LLMs) show promise for generating effective learning resources at scale for diverse groups of learners. This goes beyond performance outcomes to other intrinsic factors such as interest and motivation in learning. For example, in a between-subjects study comparing pre-existing text to AI-generated output, Leong et al. show that while there was no difference in learning performance, AI-generated output positively affected learning motivation thereby facilitating positive downstream effects (including appreciating the relevance and utility of learning the material) (Leong2024).
@@
Research in education also shows that learners benefit from having information presented in multiple modalities or formats (mayer2002). By presenting information in multimodal learning environments paired with guided activities, reflection, and feedback, learners benefit from numerous pedagogical advantages, including the promotion of independent learning and practice that is tailored to the needs of individual students (Moreno2007).
@@
The process of transforming traditional textbooks and learning materials into these more dynamic, engaging, and personalized sources can, however, be time consuming and labor intensive for teachers, thereby limiting the broader use of AI-powered education tools (dennison2025). When considering direct and independent use by students for more personalized experiences at scale, several areas of concern emerge that demand careful consideration. These include the critical issues of accuracy, bias, academic integrity, factuality and reliability of model outputs, the potential for over-reliance on AI tools and its long-term impact on cognitive abilities and critical thinking skills, and students’ inherent capacity to formulate effective prompts for facilitating their learning experience (han2024; Park2024; Tankelevitch2025; Chan2023; denny2023). Furthermore, research has been mixed when it comes to demonstrating the efficacy of AI tools on performance outcomes (do2025; Leong2024; Pallant2025).
@@
In this paper, we build upon pedagogical best practices to experimentally test the impact of an AI-powered interactive learning platform on student understanding and recall of information. We investigate the following research questions:
@@
  • RQ1 How does Learn Your Way, an experimental research platform that leverages generative AI to transform content (LYWtechreport), impact student comprehension and immediate recall of information?
@@
  • RQ2 How does Learn Your Way impact retention and recall of information over time?
@@
  • RQ3 Can Learn Your Way facilitate more engaging learning experiences for students?
@@
The results of the experiment suggest that AI-powered interactive learning platforms that incorporate content transformations and assessments, can facilitate positive learning outcomes and provide students with more engaging learning experiences compared to learning via traditional textbook chapters. We also demonstrate the positive role of timely feedback (as opposed to solutions) in equipping learners with knowledge and skills for self efficacy and academic independence.
3 Background and Related Work
3.1 HCI and Education
@@
The Human-Computer Interaction (HCI) community has a rich history of research exploring the design, development, and evaluation of educational technology (stasko1993; russell1993; abowd1998; scaife1997; zuckerman2005; Saerbeck2010; winkler2020). The discourse about generative AI in education has shifted from an initial period of denial and uncertainty to partial acceptance and explorations, ultimately leading to calls for further research (Pallant2025) and governance (unicef2019ai). Recent HCI research has explored a wide range of important topics in this domain, including privacy and security considerations for technology use in schools (kumar2019; Chanenson2023), extending participatory and co-design methods to the development of classroom-based technologies (chang2024; nicholson2022), and designing tactile-based learning experiences to foster accessibility in education (Melfi2020).
@@
Additionally, research in HCI is beginning to explore how the increase in AI-powered capabilities can be incorporated into a variety of interactive experiences in order to positively impact learning. For example, Liu et al. created and evaluated an AI agent designed to interact with students in VR-based learning environments and promote classroom participation and more positive learning experiences, finding that these agents can provide pedagogical value through increased levels of engagement and the potential for learning gains (Liu2024). Chen et al. found that augmenting block-based programing software (Scratch) with an LLM-based chat experience allowed students to learn more independently, and create higher quality projects (Chen2024). This growing body of research suggests a strong desire from the HCI community to better understand how generative AI tools can help improve student outcomes.
3.2 AI Assessment and Feedback
@@
One of the ways AI can help bring value to students is through AI-driven assessment and feedback. Pedagogical research shows that effective feedback is timely, actionable, grounded in standards or learning goals, and cites specific observations to facilitate learner understanding and support action (wiggins2012seven). Providing feedback is an inherent part of a teacher’s role, but as students partake in independent practice outside of classroom hours, feedback is not always readily available or accessible. However, the ability to access feedback is necessary to help students understand their performance and progress in relation to required standards and learning objectives (hattie2007).
@@
From the early days of intelligent tutoring systems, AI-driven knowledge assessments – including personalized, adaptive feedback – has been a core area of research for education-focused AI (anderson1985; corbett1997; graesser2004autotutor). With generative AI, students have the potential to access ongoing, targeted, timely, and individualized assessments and feedback at scale for self-directed learning and self-efficacy. Recent technological advances show improvements with the capabilities surrounding assessment and feedback (Choi2020; yang2022), but more improvements are needed to truly reach this vision (tu2025empowering).
@@
In HCI, several qualitative studies with parents, teachers, and students have led to design recommendations that emphasize the value, importance, and opportunity for AI to provide students with high-quality assessments and feedback (han2024; prasad2025). A recent meta-analysis on emerging technologies in education also emphasizes the importance of HCI research that is grounded in pedagogical strategies and focuses on assessment and feedback (vanmechelen2023). However, there is limited quantitative data and experimental studies evaluating how the design of AI-powered education tools can facilitate effective assessments and feedback. In this paper, we build upon prior work by incorporating design recommendations from past research into an experimental generative AI platform, Learn Your Way, and subsequently assessing the platform’s impact on student learning experiences and performance outcomes.
3.3 AI-Generated Content Transformations
@@
By definition, content transformations transform static content into multiple representations to help facilitate learning, such as transforming text-based content into something that is audio- and/or visual-based (AINSWORTH1999). Content transformations are an integrated process that isn’t just changing the modality (i.e., text to audio), but also the structural reorganization of information that bridges the gap between raw information and a learner’s cognitive architecture (shulman1987knowledge; Koedinger2015). In contemporary educational technology and HCI research, there is a notable rise in the adoption of generative content transformation capabilities; these extend beyond simple modality shifts to include the structural transmutation of static information into interactive assessment tools and adaptive feedback mechanisms (brusilovsky2001adaptive; mitrovic2012fifteen). This trend reflects a broader shift towards dynamic, personalized learning resources that can be adapted to various student needs and learning preferences. The benefits of content transformations are guided by principles from Mayer’s cognitive theory of multimedia learning, which states that people learn more deeply when information is received in multiple modalities as opposed to reading or hearing words alone (mayer2002; mayer1998cognitive). For example, a study showed that children who listened to stories while also viewing relevant pictures recalled more information than those who only heard stories (levin1980children). This study (and others, e.g., (Villegas2024)) suggests that integrating materials across different modalities, such as audio paired with images or videos, can be particularly beneficial for student engagement and learning.
@@
Generative AI capabilities, by their nature, create content transformations (text-to-text, text-to-audio, text-to-image, and text-to-video). By providing multiple modes of interaction and engagement with content, generative AI tools are well positioned to support self-directed learning, including theories that advocate for learners to play a central role in their own learning journeys, which has been shown to have positive impacts on learning outcomes (Deci2000; Zimmerman2009; zimmerman2000). For example, this could include metacognitive strategies that allow for setting individual learning goals, choosing appropriate strategies to achieve these goals, and monitoring one’s individual learning process and outcomes for adjustment and improvements (Panadero2017).
@@
In HCI, the research community is beginning to explore the effects of AI-based content transformations on learning and education. In an experiment, Leong et al. found that a generative AI-based vocabulary app, that transformed standard vocabulary sentences into either personalized, generated sentences or personalized, generated text-based stories did not increase learning performance but did positively impact learning motivation (Leong2024). Similarly, in an experiment moving beyond text-to-text transformations, Do et al. found that transforming educational materials from text to podcast form led to a more enjoyable learning experience and also, significantly improved learning outcomes for certain subjects (do2025).
@@
4 Learn Your Way
Learn Your Way is an experimental platform that leverages generative AI to transform textbook PDFs into a variety of formats, or multiple representations (LYWtechreport). We used Learn Your Way to test our research hypotheses. Learn Your Way is not intended to be an educational product used in classrooms. Rather, it is a way for us to test how AI and pedagogical theories have the potential of impacting learning outcomes. These content transformations were grounded in pedagogical theories that demonstrate that multiple representations of content can be a powerful way to learn (Mayer1990; mayer1998cognitive). Learn Your Way also incorporates learner control over format selection, a feature theoretically grounded in Self-Determination Theory (SDT) (ryan2000self). By enabling this choice, we aim to augment the students’ perceived autonomy, a known precursor to enhancing intrinsic motivation, cognitive effort allocation, task performance, and self-efficacy (or perceived competence) (Reber2009; SCHNEIDER2018; patall2008effects).
Learn Your Way presents five learning modalities: Immersive Text, Slides, Video, Audio Lesson, and Mindmap (Figure 1). See Supplementary Materials for a visual example of each modality.
  • Immersive Text: Transforms the original content into smaller sections releveled to a selected grade, and provides opportunities for interactive learning and practice (e.g. memory aids, quizzes, embedded questions, and timelines). These transformations of chunking and scaffolding are not just “extra features,” but mechanisms by which AI generates “dynamic” content rather than “static.” Within Immersive Text, students had access to the following features:
@@ 4 Learn Your Way
Figure 1: Diagram of the Learn Your Way user interface for Immersive Text. Other learning modalities are featured at the top including Slides, Video, Audio Lesson, and Mindmap, as well as the original PDF.
When students open Learn Your Way, they start by selecting their grade level and interests (see Appendix A). Learn Your Way then personalizes and adapts the content being generated by adjusting for grade level and incorporating the student’s selected interests. As students use Learn Your Way, they engage with interactive quizzes designed to promote active learning and assess real-time understanding of the source material, enabling further elements of personalization. Students receive dynamic feedback on their answers, can track their overall progress, and are guided to review specific content areas where they demonstrated gaps in understanding. Students are also able to view the original content, in PDF form, at any time.
4.1 Learn Your Way Design and Evaluation
Throughout the design of Learn Your Way, each modality was evaluated by pedagogical experts to ensure the fidelity and quality of each content transformation while preserving the integrity of the source content. A unified rubric was developed to assess key pedagogical criteria, including accuracy (the extent to which the AI faithfully represented the source material), coverage (the thoroughness with which all key concepts were addressed), and learning science principles. The results indicated a high level of quality, with all modalities and features achieving an average expert rating of 0.9 or higher on a scale of 0 to 1 across all pedagogical criteria. Detailed information about the rubric and results of the evaluation can be found in Supplementary Materials.
5 Experimental Study
We conducted an experimental study to compare performance outcomes and learning experiences between the AI-powered learning platform (Learn Your Way) and a digital reading of an original textbook chapter (Textbook) as a baseline comparison, consistent with established baseline comparisons employed in the literature (do2025).
The source content used in the experiment was a neuroscience textbook chapter titled, “Brain Development During Adolescence” (lumen_brain_adolescence). The source material was chosen by a panel of pedagogy experts using a set of criteria based on education psychology (see Supplementary Materials).
5.1 Research Hypotheses